Circular No. 8385

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dj IN NGC 2403 It has been noted by several contributors that star 96 of Sandage (1984,

A.J. 89, 630; B = 18.55, V = 18.31) is close to that the position of SN 2004dj (cf. IAUC 8377, 8379). H. Yamaoka, Kyushu University, remarks that star 96 was questioned by Sandage and resolved by Larsen (1999, A.Ap. Suppl. 139, 393) to be a young massive cluster. He adds that, at the distance of NGC 2403, 1′′ translates to 15 pc (as large as the scale of young clusters), and Larsen gave the effective radius of this cluster (n2403-2866) as 4.3 pc, so that the nature of this source is necessarily uncertain.

J. Maíz-Apellániz, Space Telescope Science Institute (STSI), writes that Sandage’s star 96 appears to be pointlike in the Hα image (not continuum-subtracted) obtained by Matonick et al. (1997, Ap.J. Suppl. 113, 333, Fig. 5a) but was not identified as an H II region by Sivan et al. (1990, A.Ap. 237, 23), indicating that Matonick et al. simply detected the Rband continuum. Based on these data and on a distance to NGC 2403 of 3.30 Mpc (Karachentsev et al. 2004, A.J. 127, 2031), star 96 appears to be a very bright (M_{V} < −9.3) blue supergiant. Alternatively, the progenitor could have been a red supergiant, in which case it must have been a member of a multiple system or a very compact cluster that included very bright blue stars. If the progenitor is indeed a late-B/early-A supergiant, the color (B −V = +0.24) measured by Sandage would be compatible with the value E(B − V) ∼ 0.18 measured by Patat et al. (IAUC 8378).

H. E. Bond and M. H. Siegel, STSI, report that the site of SN 2004dj was imaged on three nights in 1999 January with the Kitt Peak 4-m telescope (+ Mosaic camera) and on six nights in 1998–1999 with the Kitt Peak 0.9-m telescope. Recent CCD frames showing the supernova (provided by Y. Lipkin, D. Maoz, E. Ofek, and D. Poznanski of Wise Observatory; R. Arbour of South Wonston, Hampshire, England; D. Bishop of Hilton, NY; and O. Nielsen of Rijswijk, The Netherlands), when compared with 1999 4-m-telescope frames, indicate that SN 2004dj is within 0′′.3 of Sandage’s star 96, and no other star visible in the 4-m-telescope frames is as close to the supernova position. Star 96 is one of the brightest blue supergiants known in NGC 2403; the 0.9-m-telescope CCD frames yield V = 17.93 and B–V = +0.17. Humphreys and Aaronson (1987, A.J. 94, 1156) obtained spectra of star 96 and classified it as having a composite spectrum with an integrated early-A type but with He I lines, suggesting the presence of some B stars, as well; the object appears slightly nonstellar on the 4-m-telescope frames in 0′′.9 seeing.

2004 August 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green